OpenStudy (anonymous):

Consider f (x) = sqrt(1+x^2). (a) Show that for all c, |f'(c)| < 1. (b) Hence use the Mean Value Theorem to show that, when x != y, |sqrt(1+x^2)-sqrt(1+y^2)| < |x-y|

OpenStudy (anonymous):

f(x) = sqrt(1+x^2) with this equation, notice no matter the value of x, it will always be defined. the derivative is: x/sqrt(1+x^2) same equation, just now we have x in the numerator something to notice - x = 0, f(x) = 0. x = 1, f(x) = 1/sqrt2 x = 2, f(x) = 2/sqrt5 x = 3, f(x) = 3/sqrt10 notice the denominator keeps getting bigger and bigger, while the numerator climbs at a constant rate. for instance: x = 100, f(x) = 100/sqrt10001 another way to put it, is the denominator will always be larger than the numerator. therefore, it has to be less than 1, as it will always be a fraction.

OpenStudy (anonymous):

dont really know how to go about answering the second part.. i don't really remember what the mean value theorem is XD i assume its an extension of the fact that f'(x) will always be less than 1 though

OpenStudy (anonymous):

I believe the second part goes something like this: f(x) is continuous on the interval [a,b] and differentiable on the interval (a,b) for any a,b. By the MVT there exists: \[f'(c)=\frac{ f(b)-f(a) }{ b-a }=\frac{ \sqrt{1+b^2}-\sqrt{1+a^2 } }{ b-a } \] but we know that: \[|f'(c)|<1\] So, \[|\frac{ \sqrt{1+b^2}-\sqrt{1+a^2 } }{ b-a }|<1\] \[|\sqrt{1+b^2}-\sqrt{1+a^2 }|<|b-a|\] Now set b=x and a=y, \[|\sqrt{1+x^2}-\sqrt{1+y^2 }|<|x-y|\] x != y, as 0<0 DNE. Could anyone check what I have done is a mathematically correct argument? Thanks